At its peak, a tornado is 46.0 m in diameter and carries 505-km/h winds


5. 

At its peak, a tornado is 46.0 m in diameter and carries 505-km/h winds. What is its angular velocity in revolutions per second?


7. 

A single-engine helicopter has two rotors; a main rotor and a tail rotor. The main rotor has a diameter of 14.8 m and rotates at the rate of 420 rev/min while the tail rotor with a diameter of 1.64 m rotates at 3775 rev/min. What are the speeds, in m/s, of the tips of each rotor? 
Answer #1

5.

We know that angular speed is

=v/r

So, first of all lets make the units to be in standard system

r=23 m

And,

v=505 km/h or,

=(505*1000)/(3600) m/s

=140.278 m/s

So,

Angular speed will be

=140.278/23

=6.099 rad/sec

Now we know that

1 rad/sec= 0.159 rev/sec

So,

=6.099*0.159

=0.97 rev/sec